elimination of multiplication 7x+4y=-4
5x+8y=28

Answer by DrBeeee(684)

Given:
(1) 7x + 4y = -4 and
(2) 5x + 8y = 28
I'm not exactly sure what "elimination by multiplication" means, but I interpert it as meaning that you multiply one of the equations by by a number that allows you to eliminate one of the variables after you add or subtract the new equation from the other. If this method is what is wanted, let's multiply (1) by -2 and get
(3) -14x - 8y = 8
Now add (2) and (3) to get
(4) -9x + 0 = 36 or
(5) x = -4
Now put (5) into (1) and get
(6) 7(-4) + 4y = -4 or
(7) 4y = 24 or
(8) y = 6
Let's check the values of x and y using (2).
Is (5(-4) + 8*6 = 28)?
Is (-20 + 48 = 28)?
Is (28 = 28)? Yes
Answer: The solution pair is (-4,6)